The exchange rate (units per US dollar) for the Euro is 1.086 and for the Canadian dollar is 1.564. What is the Canadian dollar/Euro exchange rate?
Answer #1

Solution :-

Given Units per US $

So

1 US $ = Euro 1.086

Therefore 1 Euro = ( 1 / 1.086 ) US $ = 0.9208 US $

Similiarily given

1 US $ = C$ 1.564

Now We need to Calculate C$ / Euro exchange rate

C$ / Euro = ( C$ / US $ ) * ( US $ / Euro )

C$ / Euro = 1.564 * 0.9208 = 1.4401

1 Euro = C$ 1.4401

If there is any doubt please ask in comments

Thanks Please rate ....